# Robust Denoising Neural Reranker for Recommender Systems

### v1

@@ C.3 Implementation of Diverse Strategies for Leveraging Retriever Scores
Appendix D Complementary Experimental Analysis
D.1 Experiments of Retriever Stage
@@
We implement the retriever stage with the MF method [39], which aims to improve the click rate performance. The dimension of the item embedding is 128, and the learning rate is 0.001. We randomly select negative items to train the MF method, enhancing the retriever’s performance despite the item sparsity.
@@
D.2 Visualization of Learned Noises
To visualize the injected noise distribution, we compare the density curve of the injected noise from different noise injection strategies, including Gaussian noise, Beta noise, and learning-based noise. As demonstrated in Figure 5, the distribution of learned noise is more aligned with the distribution of true noise between retriever scores and user feedback. This means that the resulting posterior with model-based noise sampler better reduces the divergence term $\delta_{\text{x}}$ in Eq.(5), and potentially improve the alignment between our proposed solution with the three objectives and the data log-likelihood maximization goal.
Figure 5: The visualization of injected noise from different variants.
D.3 Detailed Results of Sensitivity Analysis
@@
D.4 Online A/B Test
Detailed Experimental Settings for Online Test The multi-stage pipeline of online system consists of three stages: retrieval, ranking, and reranking. In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage. During the ranking process, a scoring model computes multiple dimensional features for each item. Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$. DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure .
Experiment Results For our online experiments, the total traffic is randomly divided, with 12.5% allocated to the DNR method and 12.5% to the baseline method. Each experiment is conducted for at least 7 days to ensure the reliability of the results. The DNR method demonstrates a significant improvement in realshow(realshow refers to the cumulative number of videos watched by users.), increasing it by 1.089% . These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow.
Appendix E Limitation and Discussion
E.1 Limitations
@@
Finding Correct Distribution Models: In this paper, we present an implementation based on the Beta distribution assumption of the prior and the posterior. As we have mentioned in section 3.3, finding the optimal setting for the heuristic posterior might require expert experience, and we are skeptical on whether there exists a fixed setting for all user requests.
E.2 Future Directions
Multi-task Scenarios: When user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$. Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal. Thus, we consider it beneficial to further investigate how the proposed framework may accommodate the multi-task solutions.
Better Approximation for KL Divergence: In our solution, the noise regularization term $\mathcal{L}_{\text{x}}$ requires the estimation of the distribution $p_{\phi}$ and $p_{\text{x}}$. We approximate this by directly aligning the output scores with $\mathcal{L}_{\text{x}}$ in Eq.(8) in the implementation to avoid estimation of the prior. To this end, there might exist better solutions to approach the original $\mathcal{L}_{\text{x}}$ term.
More Investigation on Beta Posterior: We have illustrated that the ground truth posterior and the prior of retriever scores are likely to be a Beta distribution. Yet, there still lack analytical results on the optimal personalized parameters of $\alpha$ and $\beta$ that are aligned with the ground truth. Further study on these parameters may reveal user behavioral patterns in the real world, which could further help improve the recommendation performance.

### v2

@@ D.4 Implementation of Diverse Strategies for Leveraging Retriever Scores
Appendix E Complementary Experimental Analysis
E.1 Experiments of Retriever Stage
@@
We implement the retriever stage with the MF method Koren et al. (2009), which aims to improve the click rate performance. The dimension of the item embedding is 128, and the learning rate is 0.001. We randomly select negative items to train the MF method, enhancing the retriever’s performance despite the item sparsity.
@@
E.2 Visualization of Learned Noises
To visualize the generated noise distribution, we compare the density curve of the generated noise from different noise generators, including Gaussian noise, Beta noise, and learning-based noise. As demonstrated in Figure 5, the distribution of learned noise is more aligned with the distribution of true noise between retriever scores and user feedback. This means that the resulting posterior with model-based noise sampler better reduces the divergence term $\delta_{\text{x}}$ in Eq.(5), and potentially improve the alignment between our proposed solution with the three objectives and the data log-likelihood maximization goal.
Figure 5: The visualization of generaed noise from different variants.
E.3 Performance on Different settings of Beta-Distribution
The U-shaped Beta distribution in our paper is motivated by the binary nature of the user feedback scores and the uncertain initial guess, and it is already empirically outperforming the Gaussian distribution in our experiments. Since the heuristically designed beta distribution is not guaranteed to be optimal — unlike the model-based noise generator, whose learned noise distribution is demonstrated to be closer to the real noise distribution.
To explore different Beta distributions, we conducted additional experiments with DNR that used a beta distribution Beta(2, 5), i.e., DNR-BL, which is more aligned with the true noise distribution as well as our model-based generator (as shown in Figure 5), before switching to the model-based generator in epoch $\lambda_{e}$. Empirically, the results are presented in Table 9.
Table 9: Experiments with the beta distribution replicate the learned distribution.
Methods ML-1M Kuaivideo Book
NDCG@6 MAP@6 NDCG@6 MAP@6 NDCG@6 MAP@6
DNR-B 77.12 67.30 70.15 58.12 83.53 75.54
DNR-BL 77.43 67.49 71.10 59.61 83.68 75.79
We can observe that using this more "aligned" beta distribution (i.e., , DNR-BL) before epoch $\lambda_{e}$ offers a slight improvement compared to using the U-shaped Beta (i.e., DNR-B) before $\lambda_{e}$. This reinforces our main claim: adopting a model-based noise generator to learn and regularize noise distribution, aligning it more closely with the ground-truth noise distribution.
E.4 Detailed Results of Sensitivity Analysis
@@
E.5 Experiments on Larger Reranker Backbones
DNR is a backbone-agnostic framework for reranking tasks, utilizing and denoising the noisy scores from the previous retrieval stage through adversarial learning. Recently, LLMs have become a powerful tool for reranking tasks. However, text-based LLM4Rerank is not comparable in our setting offline or online, due to the lack of rich textual information (such as detailed item descriptions) that LLM-based rerankers fundamentally rely on. To evaluate the effectiveness of DNR in large models, we scaled up our PRM baseline by significantly increasing the number of its Transformer layers, serving as a “larger model” for comparison. We then tested our DNR framework, using this scaled-up PRM as its reranker backbone. The results are presented in the table 10.
Table 10: Experiments on larger models (expand transformers of PRM to 10 layers) with ML-1M dataset.
Methods ML-1M
HR@6 NDCG@6 MAP@6 F1@6 AUC
original PRM 60.09 72.85 62.21 65.51 88.20
original PRM+DNR 64.89 77.67 68.00 70.71 92.75
larger PRM 62.95 76.09 66.72 68.61 88.68
larger PRM+DNR 65.87 78.72 69.33 71.95 89.19
We can observe that scaling up the backbone may improve the reranking result, comparing the larger PRM with the original PRM, or comparing the larger DNR with the original DNR. Additionally, our original DNR outperforms the scaled-up “larger model” baseline (larger PRM); The DNR framework may also generalize well to a larger model.
E.6 Online A/B Test
Detailed Experimental Settings for Online Test The multi-stage pipeline of online system consists of three stages: retrieval, ranking, and reranking. In the final reranking stage, the top-6 items for user exposure are selected from a candidate set of 50 items generated by the ranking stage. During the ranking process, a scoring model computes multiple dimensional features for each item. Among these, we utilize PCTR(predicted probability that the video will be watched for more than 3 seconds) as the primary retrieval score $\mathbf{x}_{u}$, which corresponds to the “realshow” metric in the online system. Note that the final reranking stage considers the combination of the retrieval stage and the middle ranking stage as a holistic previous stage, so the output PCTR scores from the ranking stage are considered as “retrieval score” in this final stage. DNR is applied to generate a new set of scores for 50 selected items and the top-6 items are chosen as the final exposure.
Experiment Results For our online experiments, the total traffic is randomly divided, with 12.5% allocated to the DNR method and 12.5% to the baseline method. Each experiment is conducted for at least 7 days to ensure the reliability of the results. The DNR method demonstrates a significant improvement in realshow(realshow refers to the cumulative number of videos watched by users.), increasing it by 1.089% . These results indicate that the DNR method can estimate PCTR more accurately, which is strongly correlated with realshow. Additionally, the industrial recommendation environment involves various other metrics in addition to the realshow metric, including but not limited to app usage time, watch time, share-rate, like-rate, and comment-rate. We provide the corresponding results in Table 11. We can see that, in our scenario, watch-time and share-rate are slightly negatively impacted, while like-rate and comment-rate are slightly positively impacted, indicating a potential trade-off between metrics. In general, the impact on all other metrics is not statistically significant, indicating the effectiveness of our method.
metric performance boost statistically significant ($p<0.05$)
overall app-time -0.011% no
realshow +1.089% yes
watch-time -0.091% no
share-rate -0.758% no
like-rate +0.609% no
comment-rate +0.574% no
save-rate -0.690% no
profile-click-rate +0.129% no
Table 11: Extended online A/B test results.
@@ E.6 Online A/B Test
E.7 Computational Analysis
Theoretically, our method does not change the computational cost during inference since we can adopt the same reranking model structure (e.g., PRM Pei et al. (2019)). In the training phase, we include an additional noise generator that may require O(1) computational overhead, whose cost depends on the generator’s structure. Empirically, the generator does not have to be sophisticated, and the computational overhead can be trivial compared to the reranker’s training.
To better demonstrate computational cost, we compare the per-epoch training time of DNR against several baseline rerankers, including PRM Pei et al. (2019), Pier Shi et al. (2023), DCDR Lin et al. (2024), and a "w/ score" variant—where retriever scores are directly used as input to the reranker. In table 12, we present the time cost comparison for different reranker baselines in each training epoch. The variant “w/ score” denotes methods incorporating retriever scores as input features, regardless of the specific integration method (e.g., concatenation, plus, or weighted combination). Our proposed DNR demonstrates competitive training efficiency, with time costs comparable to both PRM and w/ score baselines across all datasets (ML-1M, KuaiVideo, and Book). This validates the computational efficiency of our approach while maintaining competitive performance.
Table 12: Comparison of time costs.
Datasets ML-1M Kuaivideo Book
PRM 01m01s 02m21s 01m42s
Pier 01m54s 04m11s 02m36s
DCDR 02m27s 04m58s 03m02s
w/ score 01m03s 02m49s 01m46s
DNR 01m06s 03m12s 01m51s
Appendix F Limitation and Discussion
F.1 Limitations
@@
F.2 Future Directions
Multi-task Scenarios: When a user can engage with an item with multiple behaviors (e.g., like and share a news story), DNR may adapt to this scenario by simply formulating a multi-dimensional response space for $\mathbf{z}_{u}$ and $\mathbf{x}_{u}$. Yet, different behavior signals may typically follow a heterogeneous distribution where a uniform modeling technique might become sub-optimal. Additionally, the problem may become even more sophisticated when there exist multiple retrievers that estimate the same metric, especially in industrial settings. In this case, one may have to come up with a more articulated integration method to clarify the connections between the user feedback and different retriever scores. Thus, we believe it is worth further investigating how the proposed framework may accommodate the multi-task solutions.
Multi-stage vs. End-to-End: Recent advances in end-to-end recommenders have witnessed significant advances, and evidence Deng et al. (2025) has shown that end-to-end recommenders are potentially more effective since they no longer need to worry about the consistency challenge in the multi-stage framework. Yet, there is still no evidence of a general superiority comparing these two paradigms, and there is still no optimal solution for a fully consistent multi-stage recommender.
Appendix G Use of LLMs
Large language models (LLMs) were used only to aid writing polish—including refining sentence phrasing, logical flow, and prose clarity—without altering original meanings or technical details. LLMs did not participate in core research tasks (e.g., experiment design, data processing, model training, result analysis, or drafting key technical content).